A health care professional wants to determine whether individuals with hypertension who take Atenolol have significantly lower systolic blood pressure than individuals with hypertension who do not take Atenolol. For individuals who have not been prescribed Atenolol, the population systolic blood pressure mean is 168 (µ = 168). The 30 individuals who take Atenolol have an average systolic blood pressure of 107, with a sample standard deviation of 11. On the basis of these data, can the researcher conclude that the Atenolol significantly lowers systolic blood pressure? (Use these numbers only for this question!)

Calculate the 95% confidence interval

As you know, two values will be required to complete the following equation:
  
_________ ≤ µ ≤ __________

For this question, what is the value that is below or less than µ?
Answer #1

n=30 df=n-1=29 The a54. Confidence interval for population mean() so I (triz )= 107 +(2.04x4) Lower limit= 7 - (taiz =) = 107